Evaluate an appropriate double integral (in Cartesian coordinates) to find the area of the ellipse x^2/a^2+y^2/b^2 = 1, a, b >0 What happens to your answer in the limit a→0^+? Does this make sense?
1
Expert's answer
2019-10-07T09:04:28-0400

Let us define the area "D" in Cartesian coordinates:

"y_1=-\dfrac{b}{a}\sqrt{a^2-x^2}\\
y_2=\dfrac{b}{a}\sqrt{a^2-x^2}\\
x_1=-a\\
x_2=a\\"

Then the value of the area of ellipse is given by the integral over this area "D" from the function "f(x,y)=1:\\"

"A = \iint_D \limits dxdy\\"

Opening the double integral:

"A = \int_{x_1}^{x_2}\limits dx\int_{y_1}^{y_2}\limits dy= \int_{-a}^{a}\limits dx\int_{-\frac{b}{a}\sqrt{a^2-x^2}}^{\frac{b}{a}\sqrt{a^2-x^2}}\limits dy = 2\int_{-a}^{a}\limits \frac{b}{a}\sqrt{a^2-x^2}dx=2\dfrac{b}{a}\int_{-a}^{a}\limits \sqrt{a^2-x^2}dx\\"

The last integral is the half area of circle of radius a, so it is equal to "\pi a^2/2"

"A=2\dfrac{b}{a} \dfrac{\pi a^2}{2}=\pi ab\\"

If "a" is approaching to 0, the area is approaching to 0 too.